lexxii:

In parallelogram DEFG, DH = x + 3, HF = 3y, GH = 3x – 3, and HE = 5y + 4. Find the values of x and y. x = 5, y = 12 x = 5, y = 9 x = 4, y = 9 x = 9, y = 4***** Is this right?

lexxii:

lexxii:

@Shadow

Shadow:

\[DH = HF\] \[x + 3 = 3y\] \[x = 3y - 3\] \[GH = HE\] \[3x -3 = 5y + 4\] Input x \[3(3y -3) - 3 = 5y + 4\] Solve \[9y - 9 - 3 = 5y + 4\] \[4y - 12 = 4\] \[4y = 16\] \[y =4\] Input y into first equation \[x + 3 = 3(4)\] \[x + 3 = 12\] \[x = 9\]

lexxii:

Okay I was right, thank you!

Shadow:

No problem